OpenStudy (anonymous):

An object falls freely from rest near the surface of Earth, What is the speed of the object after having fallen a distance at 4,90 meters...how do i solve this?

OpenStudy (jziggy):

Assuming zero air resistance the only force acting on the object is the force of gravity  If you use g = 9.81 and the equation: v = a*t t = 4.90 and a = 9.81 v = 4.90 * 9.81 v = 48.069 or approximately v = 48.07